OpenStudy (anonymous):

Find an equation of the given line containing the given pair of points (-3,-8) and (-9,-4)

OpenStudy (aroub):

-4+8/-9+3 4/-6 2/-3 y=mx+b, m=-2/3 , (-9,-4) -4=-2/3(-9)+b -4=6+b -10=b y=-2/3x-10

OpenStudy (anonymous):

thank u

OpenStudy (aroub):

no problem!! :D